ROLLING OF ODY ON AN INCLINED PLANE:


time taken by the body to reach the botom is given by:

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
we make the free body diagram and write equations
-ma=-mgsin(theta)+f
-I(alpha)=-Rf
alpha=a/R
fs=I*alpha/R

solving we get
a=(gsin(theta))/(1+I/MR^2)

then we find lenght pf distance to cover for it to touch ground
then 1/2*a*t^2=slant lenght